Positive charge Q is placed on a conducting spherical shell of inner radius R1 outer radius R2.A particle with charge q is placed at center of cavity.The magnitude of electric field at a point from cavity at a distance r from center is?
1
Expert's answer
2020-04-01T10:21:14-0400

According to Gauss's theorem "E \times 4 \times pi \times r^{2} = \frac{q\times r^3}{r_0\times R^3}"

"E=\frac{q\times r}{4 \pi \times R^3 \times e_0}"